# TSkips: Efficiency Through Explicit Temporal Delay Connections in Spiking Neural Networks

### v1

@@ 4.4 DVS128-Gesture, SHD and SSC Results
These improvements can be attributed to the ability of TSkips to capture long-term temporal dependencies, facilitated by the unrolled network structure, detailed in Section 3. Our method improves both the accuracy and convergence rate of the models by mitigating vanishing spikes and enabling more efficient gradient propagation during training. Analysis in the Supplementary material demonstrates faster convergence and lower energy consumption compared to standard SNNs and recurrent architectures. Additionally, we show that incorporatingTSkips in ANNs and hybrid models significantly improves gesture and speech classification.
@@ Acknowledgment
Supplementary Material
6 Inference Energy Estimation
To validate the observation that TSkips exhibit reduced energy consumption compared to baseline SNNs and standard recurrent methods (vRNNs [47] and LSTMs [19]) for various tasks (Sections 4.3 and 4.4), we estimate inference energy. These estimations are made using the method from [46, 30], which considers the energy required for sparse accumulate (AC) operations in SNNs ($E_{AC}=0.9pJ$) and dense multiply-and-accumulate (MAC) operations in ANNs ($E_{MAC}=4.6pJ$) based on a 45nm CMOS technology [20]. We calculated the total inference energy ($E_{total}$) as:
@@ 6 Inference Energy Estimation
6.1 Inference Energy Ablation Study
Figure 8: Ablation study on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on inference energy.
Building upon our ablation study in Section 4.2, we further investigated the impact of different TSkip configurations on inference energy, estimated using Eq. 6. Specifically, we explored the impact of varying temporal delay ($\Delta t$), TSkip position, and network depth on an 8-layer multilayer perceptron (MLP) baseline model with concatenation-based TSkips, found using the Neural Architecture Search (NAS) framework [27]. This analysis is crucial because TSkips increases the average spike rate within the network, potentially leading to higher inference energy consumption.
@@ 6.1 Inference Energy Ablation Study
This analysis of inference energy with varying TSkip parameters further emphasizes the complex interplay between $\Delta t$, TSkip position, and network depth. It highlights the importance of using methods like [27] to find optimal network architectures and TSkip configurations. While [27] does not explicitly optimize for minimal energy consumption, it prioritizes networks with diverse spiking activity by quantifying the dissimilarity between spiking patterns. As a result, the identified configurations often utilize TSkips with larger $\Delta t$ values, enabling our networks to efficiently capture long-term patterns while maintaining low energy consumption.
7 Additional Results - DSEC-flow
Table 4: Comparison of inference energy $(E_{total})$ using Eq. 6 for fully spiking DSEC-flow models with and without forward (F) / backward (B) TSkips across all scales.
Model #Params (×$\times 10^{6}$) $(\mathcal{N})$ Variant #OPSSNN (×$(\times 10^{9})$ Spike Rate $(\mathcal{M})$ $E_{total}$ (mJ)
@@ 7 Additional Results - DSEC-flow
F Tskip 32.7 57.8 29.4
B Tskip 30.7 54.32 27.6
@@
F Tskip 11.7 54.23 10.5
B Tskip 11.4 53.21 10.3
@@
F Tskip 6.14 65.4 5.52
B Tskip 5.86 62.4 5.27
@@
F Tskip 2.81 58.47 2.53
B Tskip 2.71 56.3 2.44
@@
F Tskip 2.29 80.6 2.07
B Tskip 2.25 78.9 2.02
7.1 Inference Energy Analysis
@@
Table 4 demonstrates the effectiveness of incorporating TSkips into fully spiking EV-FlowNet architectures [29]. By efficiently utilizing temporal information, our approach achieves comparable AEE to baseline models 3×3$\times$3 × larger, while significantly reducing inference energy. This improvement stems from TSkips ability to enhance spike propagation with minimal added parameters. While adding TSkips increases the average spike rate, the resulting models achieve comparable AEE to larger baselines (Table 1) with substantially lower energy consumption. This highlights the potential of our method for accurate, robust, and efficient optical flow prediction.
7.2 Convergence Analysis
Figure 9: Convergence analysis of fully spiking [29] and hybrid [36] EV-FlowNet architectures on the DSEC-flow dataset. The plots compare the convergence behavior of baseline models (Base and Nano) with their smaller forward (F) and backward (B) TSkip-augmented counterparts (Mini and Pico, respectively).
@@ 7.2 Convergence Analysis
As discussed in Section 3, incorporating TSkips can lead to easier training and faster convergence for both fully spiking [29]and hybrid [36] EV-FlowNet architectures. To validate this, we analyzed the convergence behavior of our DSEC-flow models across different scales. Figure 9 shows that both our largest (Base vs. Mini with TSkips) and smallest (Nano vs. Pico with TSkips) models exhibit faster convergence when incorporating TSkips, for both fully spiking and hybrid variants. This faster convergence can be attributed to TSkips ability to mitigate vanishing spikes and enhance gradient flow throughout the network, enabling quicker optimization and improved learning dynamics. Additionally, by facilitating the propagation of temporally relevant information, TSkips can help the network learn important temporal dependencies more efficiently, leading to faster convergence. This consistent trend across different model sizes highlights the benefits of TSkips in facilitating fast, efficient and effective training.
7.3 Hyper parameters for DSEC-flow
@@
Table 5 specifies the TSkip configurations used to replace existing skip connections in the fully spiking [29] and hybrid [36] EV-FlowNet architectures. Here, “TSkip Position” indicates the targeted connection, with ‘1’ representing the longest skip (between the first encoder and last decoder blocks) and ‘3’ the shortest (between the last encoder and first decoder blocks), see Fig. 5.
@@
Base + F TSkip 3 1 3 1
Base + B TSkip 4 1 5 1
Mini + F TSkip 3 1 4 2
Mini + B TSkip 3 2 3 1
Micro + F TSkip 4 2 3 2
Micro + B TSkip 3 3 3 3
Nano + F TSkip 3 1 3 2
Nano + B TSkip 4 2 3 2
Pico + F TSkip 4 1 3 1
Pico + B TSkip 3 2 4 2
8 Additional Results - DVS128 Gesture, SHD and SSC
8.1 Inference Energy Analysis
As demonstrated in Section 4.4,TSkips improves classification accuracy for speech recognition compared to standard recurrent models (vRNNs and LSTMs). Additionally, we state that TSkips also reduce inference energy compared to these models. This energy efficiency is further confirmed by the estimations presented in Table 6 for both shallow (4-layer) and deep (8-layer) MLP [47] networks on the SHD [8] dataset. These estimations were calculated using Eq. 6, which considers the energy consumed by both sparse accumulate (AC) operations in SNNs and dense multiply-and-accumulate (MAC) operations in ANNs.
@@ 8.1 Inference Energy Analysis
Model #Params (×$\times 10^{6}$) $(\mathcal{N})$OPS (×$(\times 10^{12})$ Spike Rate (×$\times 10^{4}$) $(\mathcal{M})$ $E_{\text{total}}$ (mJ)
@@
F TSkip -1 0.24 0.038 5.62 34.79
B TSkip - 1 0.19 0.022 4.95 20.37
F + B TSkip - 1 0.20 0.032 7.42 29.06
@@
F TSkip - 2 1.12 1.204 8.21 1083.90
B TSkip - 2 1.16 1.213 8.23 1091.87
F + B TSkip - 2 1.28 1.761 9.13 1585.70
8.2 Convergence Analysis

### v2

@@ 4.4 DVS128-Gesture, SHD and SSC Results
These improvements can be attributed to the ability of TSkips to capture long-term temporal dependencies, facilitated by the unrolled network structure, detailed in Section 3. Our method improves both the accuracy and convergence rate of the models by mitigating vanishing spikes and enabling more efficient gradient propagation during training. The analysis presented in Appendices A and C demonstrates lower energy consumption and faster convergence compared to standard SNNs and recurrent architectures. Additionally, we show that incorporatingTSkips in ANNs and hybrid models significantly improves speech classification (Appendix C.4).
Figure 8: Trained network’s response to temporal patterns at inference, visualized by average spiking activity. For two randomly sampled SHD event streams ((a) and (b)), we compare the true event activity against that of a baseline MLP, a forward TSkips (FTSkips) network, the FTSkips network with the skip removed (F no skip), a backward TSkips (BTSkips) network, and the BTSkips network with the skip removed (B no skip). Sub-figures (c) and (d) show the average spiking activity specifically within the forward and backward TSkips, demonstrating spike generation aligned with the specified delay window. All plots are normalized by the 700-dimensional input feature representation.
4.5 Observations
To empirically validate our claim that TSkips enhance the network’s responsiveness to temporal patterns, we analyze the spike representations of trained models at inference. Fig. 8 compares the average spiking activity of trained models with and without TSkips against the true event stream for two randomly sampled SHD data points. The TSkips models without a temporal skip have a zero tensor passing through the TSkips connection to ensure the model being evaluated does not undergo any modifications. Figures 8(a) and 8(b) show that the average spiking pattern of the trained models (even after removing the TSkips, called F/B no skip) aligns more closely with the ground truth average spike distribution compared to the baseline network. Removing TSkips during inference reduces the average spike rate and memory requirements, and the qualitative results validate that the TSkips model has learned the input event distribution and captures the input dynamics better than the baseline. We attribute this to TSkips aiding the adaptive LIF parameters and weights to learn more effectively during BPTT.
Furthermore, Fig. 8(c) and (d) isolate the average spiking activity within the forward (F) and backward (B) TSkips connections themselves. These plots reveal two key observations: first, spike generation within the TSkips aligns with the specified temporal delay $(\Delta t)$, and second, the activity mirrors a scaled version of the original input distribution. This provides strong empirical evidence that TSkips propagates relevant temporal information from the past. Furthermore, Figs. 8(a) and (b) validate that TSkips improve learning rather than acting as additional connections. The distinct characteristics of F and B TSkips are also evident. FTSkips exhibit more distributed spike patterns and higher variability, likely due to processing denser latent representations from earlier network stages. The smaller and sparser activity in BTSkips might facilitate more nuanced error correction during BPTT. Crucially, both types of TSkips demonstrate sensitivity to the entire event stream, including the sparse tail end, as seen in Figures 8(a) and (b). The sustained activity in Fig.8(a) and (b) during the later, sparser parts of the sequence in both (TSkips/no TSkips) model evaluations at inference indicates integration of information from earlier, denser time steps.
This ability of TSkips to integrate information across different temporal contexts allows the network to build richer latent representations. This mechanism is particularly important for mitigating vanishing gradients/spikes during BPTT, as TSkips establish direct pathways for information and gradients across long temporal distances, ensuring learning even when events are sparser over the sequence length. We provide quantitative results on these ‘no skip’ models for the DSEc-flow and SHD datasets in Appendixes B.2 and  C.2.
4.6 Scalability of TSkips
To assess the scalability of TSkips to deeper architectures and more complex visual tasks, we evaluated our method on the CIFAR10-DVS (Li et al., 2017) dataset. While the DSEC-flow results (Section 4.3) demonstrated the effectiveness of TSkips on a complex real-world task, the underlying EV-FlowNet architecture is custom-designed. To further validate that TSkips works on standard, widely adopted architectures in a more complex event-based setting, we applied our method to larger models like spiking ResNet18 and VGG11, incorporating the NDA data augmentation technique (Li et al., 2022b) as our baseline. The results presented in Table 4 consistently demonstrate that TSkips improve accuracy over baseline models in these deeper networks.
Table 4: Classification accuracy on the CIFAR10-DVS dataset, comparing the performance of baseline spiking ResNet18 and VGG11 models against their TSkips-augmented counterparts, illustrating the scalability of our approach. We use the NDA (Li et al., 2022b) spiking ResNet18 and VGG11 models as our baselines.
Method Base Model #Params (M) Accuracy (%)
SOTA Models
Spikingformer (Zhou et al., 2023) Single-stage ViT 2.57 81.3
$\alpha$-SSA (Xiao et al., 2025) Multi-stage ViT 1.54 82.24
SpikingResformer (Shi et al., 2024) ResNet+ViT 35.52 84.8
Basline Models
Baseline - 1 (NDA (Li et al., 2022b)) ResNet18 11.700 78.00
Baseline - 2 (NDA (Li et al., 2022b)) VGG11 132.86 81.70
TSkips Models
FTSkips - 1 ResNet18 11.708 81.08
BTSkips - 1 ResNet18 12.224 82.93
FTSkips - 2 VGG11 134.04 82.56
BTSkips - 2 VGG11 134.04 83.01
The CIFAR10-DVS dataset (Li et al., 2017) provides a valuable test case, showing that TSkips can provide accuracy gains even for event data with artificially generated temporal information from repeated closed-loop movements. These results offer insight into the effectiveness of TSkips in challenging vision/audio tasks that possess rich temporal dynamics, as detailed in Sections 4.3 and 4.4, demonstrating their ability to improve performance in deeper large-scale networks on event data. However, it is important to consider that smaller transformer models, such as Zhou et al. (2023); Xiao et al. (2025); Shi et al. (2024) can achieve similar or better performance with significantly fewer parameters.
Beyond specific task performance, the vast search space for optimal TSkips configurations (temporal delay $(\Delta t)$ and skip positions) is a recognized challenge. This motivated our use of training-free NAS for efficient exploration. As detailed in Section 3.4, the NASWOT-SAHD (Kim et al., 2022) proxy inherently guides this search by favoring networks with high linear separability in LIF activations across time steps, suggesting robustness beyond just favorable initializations. To confirm that TSkips consistently provide performance improvements over strong baselines identified by NAS, we analyzed the best-found baseline and their TSkips-augmented variants across three random seeds. These results for DSEC-flow (Gehrig et al., 2021a) are provided in Appendix D.2, Table 12, further validating the consistent improvement in inference with TSkips.
5 Discussion
While our approach demonstrates promising results, several avenues for future investigation remain. Enhancing the energy efficiency of discovered architectures is one key area, potentially through energy-based regularization within a training-free NAS proxy. Furthermore, mitigating the increase in memory usage associated with TSkips and temporal unrolling requires exploring memory-efficient training techniques or alternative architectures. Future research could also investigate making the temporal delays learnable/ dynamically changing and inference energy optimization techniques to further refine the efficiency of our method.
Implementing TSkips in neuromorphic or biological systems presents distinct challenges. The storage of historical neural states to enable connections across time steps necessitates memory allocation, which can be a significant constraint on neuromorphic hardware with limited on-chip resources. TSkips can be implemented on standard hardware by storing and retrieving past layer states, incurring memory and energy costs for simulation. The analysis in Sec. 4.5 provides a promising solution to this problem, where the trained models can be deployed without hidden state storage and passing a null input through the TSkips. On neuromorphic hardware, TSkips can be realized through direct routing with delays, where information from an earlier layer is directly routed to a later layer with a controlled temporal delay. This can be achieved using physical delay lines or digital buffers within the chip’s architecture, potentially offering energy efficiency by bypassing computations in intermediate layers. Importantly, our findings in Section 5 show that trained TSkips models can perform well even without explicitly using the temporal skip at inference, reducing hardware memory and energy requirements. Biologically, analogies might be found in short-term memory mechanisms or axonal delays that could bridge temporal gaps in information flow. Overcoming connectivity and communication latency limitations in neuromorphic hardware will also be crucial for realizing flexible temporal skips. Balancing the performance gains of TSkips with these hardware costs will guide future work exploring memory-efficient encoding of historical states, incorporating hardware constraints into the NAS process, and investigating how to mitigate the inherent hardware delays using TSkips.
@@ Acknowledgment
Appendix A Inference Energy Estimation
To validate the observation that TSkips exhibit reduced energy consumption compared to baseline SNNs and standard recurrent methods (vRNNs (Rumelhart et al., 1986) and LSTMs (Hochreiter, 1997)) for various tasks (Sections 4.3 and 4.4), we estimate inference energy. These estimations are made using the method from Rueckauer et al. (2017); Lee et al. (2021), which considers the energy required for sparse accumulate (AC) operations in SNNs ($E_{AC}=0.9pJ$) and dense multiply-and-accumulate (MAC) operations in ANNs ($E_{MAC}=4.6pJ$) based on a 45nm CMOS technology (Horowitz, 2014). We calculated the total inference energy ($E_{total}$) as:
@@ Appendix A Inference Energy Estimation
A.1 Inference Energy Ablation Study
Figure 9: Ablation study on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkips position and (c) network depth on inference energy.
Building upon our ablation study in Section 4.2, we further investigated the impact of different TSkips configurations on inference energy, estimated using Eq. 6. Specifically, we explored the impact of varying temporal delay ($\Delta t$), TSkips position, and network depth on an 8-layer multilayer perceptron (MLP) baseline model with concatenation-based TSkips, found using the Neural Architecture Search (NAS) framework (Kim et al., 2022). This analysis is crucial because TSkips increases the average spike rate within the network, potentially leading to higher inference energy consumption.
@@ A.1 Inference Energy Ablation Study
This analysis of inference energy with varying TSkips parameters further emphasizes the complex interplay between $\Delta t$, TSkips position, and network depth. It highlights the importance of using methods like Kim et al. (2022) to find optimal network architectures and TSkips configurations. While Kim et al. (2022) does not explicitly optimize for minimal energy consumption, it prioritizes networks with diverse spiking activity by quantifying the dissimilarity between spiking patterns. As a result, the identified configurations often utilize TSkips with larger $\Delta t$ values, enabling our networks to efficiently capture long-term patterns while maintaining low energy consumption.
Appendix B Additional Results - DSEC-flow
Table 5: Comparison of inference energy ($E_{total}$) using Eq. 6 for fully spiking DSEC-flow models with and without forward/backward TSkips across all scales.
Model #Params (×$\times 10^{6}$) ($\mathcal{N}$) Variant #OPSSNN (×$\times 10^{9}$) Spike Rate ($\mathcal{M}$) $E_{total}$ (mJ)
@@ Appendix B Additional Results - DSEC-flow
Forward TSkips 32.7 57.8 29.4
Backward TSkips 30.7 54.32 27.6
@@
Forward TSkips 11.7 54.23 10.5
Backward TSkips 11.4 53.21 10.3
@@
Forward TSkips 6.14 65.4 5.52
Backward TSkips 5.86 62.4 5.27
@@
Forward TSkips 2.81 58.47 2.53
Backward TSkips 2.71 56.3 2.44
@@
Forward TSkips 2.29 80.6 2.07
Backward TSkips 2.25 78.9 2.02
B.1 Inference Energy Analysis
@@
Table 5 demonstrates the effectiveness of incorporating TSkips into fully spiking EV-FlowNet architectures (Kosta & Roy, 2023). By efficiently utilizing temporal information, our approach achieves comparable AEE to baseline models 3×3$\times$3 × larger, while significantly reducing inference energy. This improvement stems from TSkips ability to enhance spike propagation with minimal added parameters. While adding TSkips increases the average spike rate, the resulting models achieve comparable AEE to larger baselines (Table 1) with substantially lower energy consumption. This highlights the potential of our method for accurate, robust, and efficient optical flow prediction.
B.2 Inference Accuracy Analysis
As detailed in Sec. 4.5 the trained TSkips models were evaluated at inference after removing the temporal hidden state propagation through the TSkips. We provide qualitative results for these models in Table 6 and see that the DSEC models retain performance even without explicitly storing temporal states at inference. We attribute this to the learnable scaling factor $(\alpha)$ within the TSkips as detailed in  3.3. This $(\alpha)$ dynamically determines the weighted contribution of both the current timestep’s information and the information from the delayed past timestep propagating through the skip connection. For our "no TSkips" inference analysis on DSEC, we used the learned $\alpha$ to control how much of the past hidden state’s contribution was zeroed out in the skip connection, allowing the model to still leverage the learned weighting of the current timestep’s information.
Table 6: Comparison of AEE (lower is better) on DSEC-flow across baseline architectures, models with forward (F)/ backward TSkips and models no TSkips.
Arch Baseline FTSkips F no TSkips BTSkips B no TSkips
Base 1.35 1.12 1.13 1.13 1.13
Micro 1.65 1.44 1.44 1.46 1.47
Mini 1.80 1.57 1.58 1.56 1.56
Nano 2.17 1.86 1.86 1.77 1.77
Pico 2.57 2.19 2.20 2.28 2.30
B.3 Convergence Analysis
@@
As discussed in Section 3, incorporating TSkips can lead to easier training and faster convergence for both fully spiking (Kosta & Roy, 2023)and hybrid (Negi et al., 2024) EV-FlowNet architectures. To validate this, we analyzed the convergence behavior of our DSEC-flow models across different scales. Fig. 10 shows that both our largest (Base vs. Mini with TSkips) and smallest (Nano vs. Pico with TSkips) models exhibit faster convergence when incorporating TSkips, for both fully spiking and hybrid variants. This faster convergence can be attributed to TSkips ability to mitigate vanishing spikes and enhance gradient flow throughout the network, enabling quicker optimization and improved learning dynamics. Additionally, by facilitating the propagation of temporally relevant information, TSkips can help the network learn important temporal dependencies more efficiently, leading to faster convergence. This consistent trend across different model sizes highlights the benefits of TSkips in facilitating fast, efficient and effective training.
Figure 10: Convergence analysis of fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architectures on the DSEC-flow dataset. The plots compare the convergence behavior of baseline models (Base and Nano) with their smaller forward (F) and backward (B) TSkips -augmented counterparts (Mini and Pico, respectively).
B.4 Hyper parameters for DSEC-flow
@@
Table 7 specifies the TSkips configurations used to replace existing skip connections in the fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architectures. Here, “TSkips Position" indicates the targeted connection, with ‘1’ representing the longest skip (between the first encoder and last decoder blocks) and ‘3’ the shortest (between the last encoder and first decoder blocks), see Fig. 5.
@@
Base + Forward TSkips 3 1 3 1
Base + Backward TSkips 4 1 5 1
Mini + Forward TSkips 3 1 4 2
Mini + Backward TSkips 3 2 3 1
Micro + Forward TSkips 4 2 3 2
Micro + Backward TSkips 3 3 3 3
Nano + Forward TSkips 3 1 3 2
Nano + Backward TSkips 4 2 3 2
Pico + Forward TSkips 4 1 3 1
Pico + Backward TSkips 3 2 4 2
Appendix C Additional Results - DVS128 Gesture, SHD and SSC
C.1 Inference Energy Analysis
As demonstrated in Section 4.4,TSkips improves classification accuracy for speech recognition compared to standard recurrent models (vRNNs and LSTMs). Additionally, we state that TSkips also reduce inference energy compared to these models. This energy efficiency is further confirmed by the estimations presented in Table 8 for both shallow (4-layer) and deep (8-layer) MLP (Rumelhart et al., 1986) networks on the SHD (Cramer et al., 2022) dataset. These estimations were calculated using Eq. 6, which considers the energy consumed by both sparse accumulate (AC) operations in SNNs and dense multiply-and-accumulate (MAC) operations in ANNs.
@@ C.1 Inference Energy Analysis
Model
#Params
(×$\times 10^{6}$)
($\mathcal{N}$)
#OPS
(×$\times 10^{12}$)
Spike Rate
(×$\times 10^{4}$)
($\mathcal{M}$)
$E_{\text{total}}$
@@
Forward TSkips -1 0.24 0.038 5.62 34.79
Backward TSkips - 1 0.19 0.022 4.95 20.37
Forward + Backward TSkips - 1 0.20 0.032 7.42 29.06
@@
Forward TSkips - 2 1.12 1.204 8.21 1083.90
Backward TSkips - 2 1.16 1.213 8.23 1091.87
Forward + Backward TSkips - 2 1.28 1.761 9.13 1585.70
C.2 Inference Accuracy Analysis
Table 9: Comparison of accuracy on SHD for two baseline models, forward (F)/ backward TSkips and no TSkips.
Model TSkips no TSkips
Baseline - 1 84.32
FTSkips - 1 92.32 90.71
BTSkips - 1 93.64 92.78
F+B TSkips - 1 93.01 91.76
Baseline - 2 86.42
FTSkips - 2 94.15 92.32
BTSkips - 2 93.86 93.18
F+B TSkips - 2 94.73 92.34
As detailed in Sec. 4.5 the trained TSkips models were evaluated at inference after removing the temporal hidden state propagation through the TSkips. We see that the SHD models without the TSkips have a slight decrease in inference accuracy. We attribute this slight decrease to less information being propagated to the later layers of the network through the TSkips. However, the models continue to perform better than the baselines, thus validating that TSkips helped to improve learning during training.
C.3 Convergence Analysis